# Quantum machines using $\mathrm{Cu}_{3}$-like compounds modeled by Heisenberg antiferromagnetic in a triangular ring

### v2

@@ II Thermodynamics of the model
To investigate quantum thermodynamics for machine applications, we analyze the energy levels of a spin Hamiltonian relevant to ${\rm Cu}_{3}$-like systems, using the Heisenberg model in an isosceles triangular spin ring context[36, 38, 39]. This study builds on previous work that characterizes the Hamiltonian of ${\rm Cu}_{3}$ compounds [36, 38, 39]. The Hamiltonian is expressed as
@@
where $J_{j,j+1}^{\alpha}$ represents the exchange interaction parameters between sites $j$ and $j+1$ for $\alpha=\{x,y,z\}$. The vector ${\bf D}_{j,j+1}=(D_{j,j+1}^{x},D_{j,j+1}^{y},D_{j,j+1}^{z})$ represents the Dzyaloshinskii-Moriya interaction, while $\mathbf{g}_{j}=(g_{j}^{x},g_{j}^{y},g_{j}^{z})$ are the site-dependent g-factors. The magnetic field $\mathbf{B}$ is assumed to be site-independent on the triangle, and $\mu_{B}$ is the Bohr magneton. Parameters were derived from Electron Spin Resonance (ESR) data [36, 38, 39], shown in table 1. Only ${\bf D}_{1,2}=(D,D,D)$ is isotropic, while ${\bf D}_{2,3}$ and ${\bf D}_{3,1}$ contribute only to the z-component, ${\bf D}_{2,3}={\bf D}_{3,1}=(0,0,D)$.
@@
Table 1: Magnetic parameters of the ${\rm Cu}_{3}-X$} compounds adjusted to experimental data in units of Kelvin, where $X$ denotes either As or Sb, as extracted from reference Magnetic[38].
@@
where the eigenvalues $\varepsilon_{i}(B)$ depend on the Hamiltonian parameters (table 1), the magnetic field $\mathbf{B}$ (in Tesla), and the temperature $T$ (in Kelvin). Although any physical quantity can be derived from $\mathcal{Z}$, the eigenvalues must be obtained numerically, limiting analytical solutions. This result, relevant to quantum machines, highlights the importance of understanding magnetic and thermodynamic properties at the quantum level, crucial for advancing quantum technologies.
@@ II.3 Thermal efficiency
This new metric evaluates COP within a bounded range of $10<\kappa<$ 1, implying that as COP approaches 00, $\kappa$ similarly trends toward 00, while $\kappa$ approaches $1$ as COP tends to infinity. Additionally, a $COP=1$ corresponds to $\kappa=0.5$. Consequently, the thermal efficiency for all operational modes can be defined within the range of 00 to $1$, with 00 indicating the worst performance and $1$ the best. This definition is applicable for the refrigerator, heater, and accelerator, each described in Table 2.
For our discussion, we will focus on the ${\rm Cu}_{3}-{\rm As}$ compound. The ${\rm Cu}_{3}-{\rm Sb}$ compound demonstrates similar characteristics, as shown by the comparable parameters in Table 1.
@@ III Quantum Carnot machine
Figure 2: Operational modes of the quantum Carnot cycle for ${\rm Cu}_{3}$-like compound in the $B_{0}$-$B_{1}$ plane. The green region operates as engine, while the cyan region operates as refrigerator. (a) For $T_{l}=0.5$K and $T_{h}=1$K. (b) For $T_{l}=0.7$K and $T_{h}=1$K. (c) For $T_{l}=0.7$K and $T_{h}=1.5$K. (d) For $T_{l}=1$K and $T_{h}=1.5$K.
In Fig. 2, we show the operational modes of the quantum Carnot cycle for a ${\rm Cu}_{3}$-like compound in the $B_{0}$ vs. $B_{1}$ plane. The green region indicates the heat engine mode, and the cyan region indicates the refrigerator mode. Panel (a) displays for cold reservoir $T_{l}=0.5$K and hot reservoir $T_{h}=1$K. Assuming $B_{0}=0$, the system operates as a heat engine for $B_{1}>1.2$ and as a refrigerator at $B_{1}\approx 4.7$T, where MCE is enhanced [52]. The heat engine efficiency is $\eta=0.5$ in the green region, with a refrigerator $COP$ of 2. Panel (b) for $T_{l}=0.7$K and $T_{h}=1$K shows similar behavior, with refrigeration around $B_{1}\approx 4.7$T and a small heat engine region at $B_{0}\approx 5$T and $B_{1}\approx 3$T. The efficiency is $\eta=0.3$, with a COP of 3.333. In panel (c) with $(T_{l}=0.7$K, $T_{h}=1.5$K), the small heat engine region disappears, but operation continues at $B_{0}\approx 0T$ and $B_{1}\approx 4.7$T. The efficiency is $\eta=0.533$, with a COP of 1.875. Panel (d) with $T_{l}=1$K and $T_{h}=1.5$K resembles panel (a), with heat engine operation for $B_{1}>2.5$T, efficiency $\eta=0.333$, while the COP for the refrigerator is $COP=3$.
@@
Figure 3: Operational modes of the quantum Carnot cycle for ${\rm Cu}_{3}$-like compound in the $B_{1}-T_{h}$ plane. Below the pink curve operates as a heat engine, while above the curve operates as a refrigerator. (a) For $T_{l}=0.1$K and $B_{0}=0$T. (b) For $T_{l}=0.5$K and $B_{0}=0$T.
@@ IV Quantum Otto machine
Now we will look at how a compound similar to ${\rm Cu}_{3}$ could work as a quantum Otto machine. This study is important for the field of quantum thermodynamics, as it offers insights into how ${\rm Cu}_{3}$-like quantum systems might be used for energy conversion. Understanding the quantum versions of classical thermodynamic processes can help us better manage energy at a tiny, microscopic scale.
@@
    Isomagnetic process (b-c): In this stage, the system comes into thermal contact with a hot reservoir, and the temperature increases while keeping the magnetic field constant. No work is done,$W_{bc}=0$. Consequently, the system absorbs heat, which is given by
@@
    Isomagnetic process (d-a): In this process, the system comes into contact with a cold environment, gradually lowering its temperature. Since there is no change in the magnetic field, no work is done ($W_{da}=0$). The heat released to the environment is given by
@@
Figure 5: Operational modes of the quantum Otto cycle for ${\rm Cu}_{3}$-like compound in the $B_{0}-B_{1}$ plane. In green region operates as engine, in cyan region operates as refrigerator, the yellow region operates as heater, and in red region operates accelerator. (a) For $T_{l}=0.5$K and $T_{h}=1$K. (b) For $T_{l}=0.7$K and $T_{h}=1$K. (c) For $T_{l}=0.7$K and $T_{h}=1.5$K. (d) For $T_{l}=1$K and $T_{h}=1.5$K.
In Fig. 5, we present the operational modes of the quantum Otto cycle for a ${\rm Cu}_{3}$-like compound in the $B_{0}-B_{1}$ plane. The color-coded regions signify different modes of operation: the green region represents the engine mode, the cyan region represents the refrigerator mode, the yellow region corresponds to the heater mode, and the red region indicates the accelerator mode. In Panel (a), we consider the case with $T_{l}=0.5$K and $T_{h}=1$K. Just to explore the MCE, we assume conveniently $B_{0}$ is approximately $0.1$T. Notably, the system exhibits peculiar behavior in the range of $2.2{\rm T}\lesssim B_{1}\lesssim 3.8{\rm T}$, where it functions as a refrigerator. However, at $B_{1}\approx 4$T, where the MCE is significantly enhanced, the system operates as a thermal accelerator. For $B_{0}>B_{1}$, the system predominantly acts as a thermal accelerator, while for $B_{1}>B_{0}$, it exhibits four operation modes. Specifically, for $B_{1}\gtrsim B_{0}$, the system serves as a heat engine, and where the MCE is enhanced, it behaves as an accelerator. In Panel (b), with $T_{l}=0.7$K and $T_{h}=1$K, we observe similar behavior. Although in this case, the ${\rm Cu}_{3}$-like compound could work as a refrigerator at $B_{1}\approx 1$T. However, a small region appears as a refrigerator at $B_{0}\approx 5$T, surrounded by a heater operation mode. In Panel (c), for $T_{l}=0.7$K and $T_{h}=1.5$K, the region where the system operates as a refrigerator and heater at $B_{0}\approx 5$T and $B_{1}\approx 3$T disappears. However, for $B_{0}\approx 0.1$T and $B_{1}\approx 3$T, the system continues to operate as a refrigerator. Similarly, in Panel (d), for $T_{l}=1$K and $T_{h}=1.5$K, the behavior is somewhat akin to panel (a), although for $B_{0}=0$ and $B_{1}\gtrsim 5$T, it operates as a refrigerator.
@@
Figure 6: Thermal efficiency of the quantum Otto cycle for ${\rm Cu}_{3}$-like compound in the $B_{0}$-$B_{1}$ plane. The cycle is evaluated under varying cold ($T_{l}$) and hot ($T_{h}$) temperature conditions: (a) For $T_{l}=0.5$ and $T_{h}=1$. (b) For $T_{l}=0.7$ and $T_{h}=1$. (c) For $T_{l}=0.7$ and $T_{h}=1.5$. (d) For $T_{l}=1$ and $T_{h}=1.5$.
@@
Additionally, we can analyze the quantum operational modes as functions of the hot temperature $T_{h}$ and magnetic field $B_{1}$, assuming $B_{0}=0$T. In contrast to the Carnot engine, thermal efficiency now depends on the magnetic fields $B_{0}$ and $B_{1}$, as well as the temperatures $T_{l}$ and $T_{h}$. In Fig.7, Panel (a) illustrates the operational modes of the quantum Otto cycle for a ${\rm Cu}_{3}$-like compound in the $B_{1}-T_{h}$ plane, assuming fixed values of $T_{l}=0.1$K and $B_{0}=0$. Under these conditions, we observe all four operational modes: engine (green), refrigerator (cyan), heater (yellow), and accelerator (red). Panel (b) presents a similar plot but assumes $T_{l}=0.5$K and $B_{0}=0$T. It demonstrates that at lower temperatures, the system can function as a refrigerator, although this region significantly decreases for lower $T_{h}$ values. Conversely, for higher temperatures, the system tends to operate as a heat engine. In panels (c-d), we depict the thermal efficiency under the same conditions as in panels (a-b), respectively. For thermal efficiency, we use $\eta$ for the heat engine and $\kappa$ for the refrigerator, heater, and accelerator, as detailed in Table 2. Here, we can observe that thermal efficiency depends on $T_{h}$ and $B_{1}$. The COP of the refrigerator is roughly around $COP=1$ or $\kappa=0.5$, while the efficiency of the heat engine is approximately $\eta\sim 0.2$. The COP for the heater is also quite small. However, the efficiency of the thermal accelerator is notably strong, approaching $\kappa\rightarrow 1$.
Figure 7: (a) Operation mode of the quantum Otto cycle for ${\rm Cu}_{3}$-like compound in the plane $B_{1}-T_{h}$, for $T_{l}=0.1$ and $B_{0}=0$. (b) For $T_{l}=0.5$ and $B_{0}=0$. (c-d) Thermal efficiency under the same condition to (a-b) respectively.
Hence, the ${\rm Cu}_{3}$-like compound may offer certain advantages, mainly in low-temperature region around MCE is relevant, around $B_{1}\sim 4.7$T, and for temperature $T_{h}\gtrsim 2$K, the heat engine operates with high efficiency roughly around $\eta\sim 0.5$.
@@ V Quantum Stirling machine
Finally, we consider how ${\rm Cu}_{3}$-like compound could work as a quantum Stirling machine which represents a relevant stride in the field of quantum thermodynamics, offering insights into the manipulation and potential applicability of quantum ${\rm Cu}_{3}$-like systems for energy conversion. By understanding the quantum mechanical counterparts of classical thermodynamic processes, could help the energy management at microscopic scales.
@@
    and the heat released to environment is
@@
    Isomagnetic process (b-c). In this process no work is realized by the compound, because the magnetic field holds constant $W_{bc}=0$. However, the temperature of the system increases to $T_{h}$ increasing the magnetic entropy and spins alignments becoming more erratic, consequently the system absorb heat from the reservoir, which is given by
@@
    Isomagnetic process (d-a). Here the ${\rm Cu}_{3}$-like system is placed in contact with cold bath consequently decreasing the magnetic entropy and then the temperature of the system also decreases, but since there is no change of magnetic field, then no work is realized $W_{da}=0$. Thus the heat released to the cold bath by the system is given by
@@
Figure 9: Operational modes of the quantum Stirling cycle for ${\rm Cu}_{3}$-like compound in the $B_{0}$-$B_{1}$ plane. In green region operates as engine, in cyan region operates as refrigerator, the yellow region operates as heater, and in red region operates accelerator. (a) For $T_{l}=0.5$K and $T_{h}=1$K. (b) For $T_{l}=0.7$K and $T_{h}=1$K. (c) For $T_{l}=0.7$K and $T_{h}=1.5$K. (d) For $T_{l}=1$K and $T_{h}=1.5$K.
In Fig.9, we present the quantum operational modes of the Stirling cycle for an engine using a ${\rm Cu}_{3}$-like compound in the $B_{0}-B_{1}$ plane. Panel (a) illustrates this for fixed temperatures $T_{l}=0.5$K and $T_{h}=1$K. It is clear that when $B_{1}>B_{0}$, the system mainly operates as a heat engine. Conversely, when $B_{1}<B_{0}$, the system functions as a refrigerator. Interestingly, around $B_{0}\sim 4.5$T and $B_{1}\sim 3$T, there is a region where it operates as a heat engine. In contrast, around $B_{0}\sim 3$T and $B_{1}\sim 4.5$T, the system functions as an accelerator. There is even a small region at $B_{0}\sim 3$T and $B_{1}\sim 5$T where it operates as a heater. Additionally, between these two main operational modes, there are regions where the heater and thermal accelerator modes occur, particularly when $B_{1}\lesssim B_{0}$. Panel (b) depicts the scenario for $T_{l}=0.7$K and $T_{h}=1$K, showing a similar pattern to panel (a). The main difference is that the heater operation mode is significantly reduced. However, at $B_{0}\sim 3$T and $B_{1}\sim 5$, the system operates as a refrigerator, surrounded by the heater operation mode. Moreover, for $B_{0}\sim 4.5$T and $B_{1}\lesssim 1$T, the system acts as a refrigerator. In Panel (c), we plot the results for $T_{l}=0.7$K and $T_{h}=1.5$K, which is again similar to (b). However, there are no heater or refrigerator modes at $B_{0}\sim 3$T and $B_{1}\sim 5$, though the accelerator mode still remains at $B_{0}\sim 4.5$T and $B_{1}\lesssim 1$T. Finally, Panel (d) presents results for a different set of parameters, $T_{l}=1$K and $T_{h}=1.5$K. Here, we observe wider regions for the heater and accelerator modes, but the inversion operation mode still occurs at $B_{0}\sim 3$T and $B_{1}\sim 4.5$T, and vice versa.
@@
Figure 10: Thermal efficiency of the quantum Stirling cycle for ${\rm Cu}_{3}$-like compound in the $B_{0}$-$B_{1}$ plane. The cycle is evaluated under varying cold ($T_{l}$) and hot ($T_{h}$) temperature conditions: (a) For $T_{l}=0.5$ and $T_{h}=1$. (b) For $T_{l}=0.7$ and $T_{h}=1$. (c) For $T_{l}=0.7$ and $T_{h}=1.5$. (d) For $T_{l}=1$ and $T_{h}=1.5$.

### v3

@@ II Thermodynamics of the model
To investigate quantum thermodynamics for machine applications, we analyze the energy levels of a spin Hamiltonian relevant to ${\rm Cu}_{3}$-like systems, using the Heisenberg model in an isosceles triangular spin ring context[36, 38, 39]. This study builds on previous work that characterizes the Hamiltonian of ${\rm Cu}_{3}-X$ compounds [36, 38, 39]. The Hamiltonian is expressed as
@@
where $J_{j,j+1}^{\alpha}$ represents the exchange interaction parameters between sites $j$ and $j+1$ for $\alpha=\{x,y,z\}$. The vector ${\bf D}_{j,j+1}=(D_{j,j+1}^{x},D_{j,j+1}^{y},D_{j,j+1}^{z})$ represents the Dzyaloshinskii-Moriya interaction, while $\mathbf{g}_{j}=(g_{j}^{x},g_{j}^{y},g_{j}^{z})$ are the site-dependent g-factors. The magnetic field $\mathbf{B}$ is assumed to be site-independent on the triangle, and $\mu_{B}$ is the Bohr magneton. Parameters were derived from Electron Spin Resonance (ESR) data [36, 38, 39], shown in table 1. Only ${\bf D}_{1,2}=(D,D,D)$ is isotropic, while ${\bf D}_{2,3}$ and ${\bf D}_{3,1}$ contribute only to the z-component, ${\bf D}_{2,3}={\bf D}_{3,1}=(0,0,D)$. A schematic view of the ${\rm Cu}_{3}$-like system is illustrated in Fig.1.
Figure 1: Schematic representation of the ${\rm Cu}_{3}$-like system, where the triangle lies in the $xy$-plane.
@@
Table 1: Magnetic parameters of the ${\rm Cu}_{3}-X$} compounds adjusted to experimental data in units of Kelvin, where $X$ denotes either As or Sb, as extracted from Ref. [38].
@@
where the eigenvalues $\varepsilon_{i}(B)$ depend of the Hamiltonian parameters given in table 1, external magnetic field $\mathbf{B}$ (in Tesla), and the temperature $T$ (in Kelvin). Although any physical quantity can be derived from $\mathcal{Z}$, the eigenvalues must be obtained numerically, limiting analytical solutions. This result, relevant to quantum machines, highlights the importance of understanding magnetic and thermodynamic properties at the quantum level, crucial for advancing quantum technologies.
@@ II.3 Thermal efficiency
This new metric evaluates COP within a bounded range of $10<\kappa<$ 1, implying that as COP approaches 00, $\kappa$ similarly trends toward 00, while $\kappa$ approaches $1$ as COP tends to infinity. Additionally, a $COP=1$ corresponds to $\kappa=0.5$. Consequently, the thermal efficiency for all operational modes can be defined within the range of 00 to $1$, with 00 indicating the worst performance and $1$ the best. This definition is applicable to the refrigerator, heater, and accelerator, each described in Table 2.
From now on, our discussion will focus on the ${\rm Cu}_{3}-{\rm As}$ compound, as the ${\rm Cu}_{3}-{\rm Sb}$ compound demonstrates similar characteristics, with parameters values closely comparable to those of ${\rm Cu}_{3}-{\rm As}$, as shown in Table 1.
@@ III Quantum Carnot machine
Figure 3: Operational modes of the quantum Carnot cycle for ${\rm Cu}_{3}$-As compound in the $B_{0}$-$B_{1}$ plane. The green region operates as engine, while the cyan region operates as refrigerator. (a) For $T_{l}=0.5$K and $T_{h}=1$K. (b) For $T_{l}=0.7$K and $T_{h}=1$K. (c) For $T_{l}=0.7$K and $T_{h}=1.5$K. (d) For $T_{l}=1$K and $T_{h}=1.5$K.
In Fig. 3, we show the operational modes of the quantum Carnot cycle for a ${\rm Cu}_{3}$-As compound in the $B_{0}$ vs. $B_{1}$ plane. The green region indicates the heat engine mode, and the cyan region indicates the refrigerator mode. Panel (a) displays for cold reservoir $T_{l}=0.5$K and hot reservoir $T_{h}=1$K. Assuming $B_{0}=0$, the system operates as a heat engine for $B_{1}>1.2$ and as a refrigerator at $B_{1}\approx 4.7$T. The temperatures and magnetic fields values were chosen such that the MCE is enhanced[52]. The heat engine efficiency is $\eta=0.5$ in the green region, with a refrigerator $COP$ of 2. Panel (b) for $T_{l}=0.7$K and $T_{h}=1$K shows similar behavior, with refrigeration around $B_{1}\approx 4.7$T and a small heat engine region at $B_{0}\approx 5$T and $B_{1}\approx 3$T. The efficiency is $\eta=0.3$, with a COP of 3.333. In panel (c) with $(T_{l}=0.7$K, $T_{h}=1.5$K), the small heat engine region disappears, but operation continues at $B_{0}\approx 0T$ and $B_{1}\approx 4.7$T. The efficiency is $\eta=0.533$, with a COP of 1.875. Panel (d) with $T_{l}=1$K and $T_{h}=1.5$K resembles panel (a), with heat engine operation for $B_{1}>2.5$T, efficiency $\eta=0.333$, while the COP for the refrigerator is $COP=3$.
@@
Figure 4: Operational modes of the quantum Carnot cycle for ${\rm Cu}_{3}$-As compound in the $B_{1}-T_{h}$ plane. Below the pink curve operates as a heat engine, while above the curve operates as a refrigerator. (a) For $T_{l}=0.1$K and $B_{0}=0$T. (b) For $T_{l}=0.5$K and $B_{0}=0$T.
@@ IV Quantum Otto machine
Now we will look at how a compound similar to ${\rm Cu}_{3}$ could work as a quantum Otto machine as illustrated in Fig.5. This study is important for the field of quantum thermodynamics, as it offers insights into how ${\rm Cu}_{3}$-like quantum systems might be used for energy conversion. Understanding the quantum versions of classical thermodynamic processes can help us better manage energy at a tiny, microscopic scale.
@@
    isochoric process (b-c): In this stage, the system comes into thermal contact with a hot reservoir, and the temperature increases while keeping the magnetic field constant. No work is done,$W_{bc}=0$. Consequently, the system absorbs heat, which is given by
@@
    isochoric process (d-a): In this process, the system comes into contact with a cold environment, gradually lowering its temperature. Since there is no change in the magnetic field, no work is done ($W_{da}=0$). The heat released to the environment is given by
@@
Figure 6: Operational modes of the quantum Otto cycle for ${\rm Cu}_{3}$-As compound in the $B_{0}-B_{1}$ plane. In green region operates as an engine, in cyan region operates as refrigerator, the yellow region operates as heater, and in red region operates accelerator. (a) For $T_{l}=0.5$K and $T_{h}=1$K. (b) For $T_{l}=0.7$K and $T_{h}=1$K. (c) For $T_{l}=0.7$K and $T_{h}=1.5$K. (d) For $T_{l}=1$K and $T_{h}=1.5$K.
In Fig. 6, we present the operational modes of the quantum Otto cycle for a ${\rm Cu}_{3}$-As compound in the $B_{0}-B_{1}$ plane. The color-coded regions signify different modes of operation: the green region represents the engine mode, the cyan region represents the refrigerator mode, the yellow region corresponds to the heater mode, and the red region indicates the accelerator mode. In Panel (a), we consider the case with $T_{l}=0.5$K and $T_{h}=1$K, the choice of these parameters is related to the MCE enhancement. Just to explore the MCE, we assume conveniently $B_{0}$ is approximately $0.1$T. Notably, the system exhibits peculiar behavior in the range of $2.2{\rm T}\lesssim B_{1}\lesssim 3.8{\rm T}$, where it functions as a refrigerator. However, at $B_{1}\approx 4$T, where the MCE is significantly enhanced, the system operates as a thermal accelerator. For $B_{0}>B_{1}$, the system predominantly acts as a thermal accelerator, while for $B_{1}>B_{0}$, it exhibits four operation modes. Specifically, for $B_{1}\gtrsim B_{0}$, the system serves as a heat engine, and where the MCE is enhanced, it behaves as an accelerator. In Panel (b), with $T_{l}=0.7$K and $T_{h}=1$K, we observe similar behavior. Although in this case, the ${\rm Cu}_{3}$-As compound could work as a refrigerator at $B_{1}\approx 1$T. However, a small region appears as a refrigerator at $B_{0}\approx 5$T, surrounded by a heater operation mode. In Panel (c), for $T_{l}=0.7$K and $T_{h}=1.5$K, the region where the system operates as a refrigerator and heater at $B_{0}\approx 5$T and $B_{1}\approx 3$T disappears. However, for $B_{0}\approx 0.1$T and $B_{1}\approx 3$T, the system continues to operate as a refrigerator. Similarly, in Panel (d), for $T_{l}=1$K and $T_{h}=1.5$K, the behavior is somewhat akin to panel (a), although for $B_{0}=0$ and $B_{1}\gtrsim 5$T, it operates as a refrigerator.
@@
Figure 7: Thermal efficiency of the quantum Otto cycle for ${\rm Cu}_{3}$-As compound in the $B_{0}$-$B_{1}$ plane. The cycle is evaluated under varying cold ($T_{l}$) and hot ($T_{h}$) temperature conditions: (a) For $T_{l}=0.5$ and $T_{h}=1$. (b) For $T_{l}=0.7$ and $T_{h}=1$. (c) For $T_{l}=0.7$ and $T_{h}=1.5$. (d) For $T_{l}=1$ and $T_{h}=1.5$.
@@
Additionally, we can analyze the quantum operational modes as functions of the hot temperature $T_{h}$ and magnetic field $B_{1}$, assuming $B_{0}=0$T. In contrast to the Carnot engine, thermal efficiency now depends on the magnetic fields $B_{0}$ and $B_{1}$, as well as the temperatures $T_{l}$ and $T_{h}$. In Fig.8, Panel (a) illustrates the operational modes of the quantum Otto cycle for a ${\rm Cu}_{3}$-As compound in the $B_{1}-T_{h}$ plane, assuming fixed values of $T_{l}=0.1$K and $B_{0}=0$. Under these conditions, we observe all four operational modes: engine (green), refrigerator (cyan), heater (yellow), and accelerator (red). Panel (b) presents a similar plot but assumes $T_{l}=0.5$K and $B_{0}=0$T. It demonstrates that at lower temperatures, the system can function as a refrigerator, although this region significantly decreases for lower $T_{h}$ values. Conversely, for higher temperatures, the system tends to operate as a heat engine. In panels (c-d), we depict the thermal efficiency under the same conditions as in panels (a-b), respectively. For thermal efficiency, we use $\eta$ for the heat engine and $\kappa$ for the refrigerator, heater, and accelerator, as detailed in Table 2. Here, we can observe that thermal efficiency depends on $T_{h}$ and $B_{1}$. The COP of the refrigerator is roughly around $COP=1$ or $\kappa=0.5$, while the efficiency of the heat engine is approximately $\eta\sim 0.2$. The COP for the heater is also quite small. However, the efficiency of the thermal accelerator is notably strong, approaching $\kappa\rightarrow 1$.
Figure 8: (a) Operation mode of the quantum Otto cycle for ${\rm Cu}_{3}$-As compound in the plane $B_{1}-T_{h}$, for $T_{l}=0.1$ and $B_{0}=0$. (b) For $T_{l}=0.5$ and $B_{0}=0$. (c-d) Thermal efficiency under the same condition to (a-b) respectively.
Hence, the ${\rm Cu}_{3}$-As compound may offer certain advantages, mainly in low-temperature region around MCE is relevant, around $B_{1}\sim 4.7$T, and for temperature $T_{h}\gtrsim 2$K, the heat engine operates with high efficiency roughly around $\eta\sim 0.5$.
@@ V Quantum Stirling machine
Finally, we consider how ${\rm Cu}_{3}$-like compound could work as a quantum Stirling machine (see Fig. 9)which represents a relevant stride in the field of quantum thermodynamics, offering insights into the manipulation and potential applicability of quantum ${\rm Cu}_{3}$-like systems for energy conversion. By understanding the quantum mechanical counterparts of classical thermodynamic processes, could help the energy management at microscopic scales.
@@
    and the heat released to the environment is
@@
    isochoric process (b-c). In this process no work is realized by the compound, because the magnetic field holds constant $W_{bc}=0$. However, the temperature of the system increases to $T_{h}$ increasing the magnetic entropy and spins alignments becoming more erratic, consequently the system absorb heat from the reservoir, which is given by
@@
    isochoric process (d-a). Here the ${\rm Cu}_{3}$-like system is placed in contact with cold bath consequently decreasing the magnetic entropy and then the temperature of the system also decreases, but since there is no change of magnetic field, then no work is realized $W_{da}=0$. Thus the heat released to the cold bath by the system is given by
@@
Figure 10: Operational modes of the quantum Stirling cycle for ${\rm Cu}_{3}$-As compound in the $B_{0}$-$B_{1}$ plane. In green region operates as an engine, in cyan region operates as a refrigerator, the yellow region operates as a heater, and in red region operates accelerator. (a) For $T_{l}=0.5$K and $T_{h}=1$K. (b) For $T_{l}=0.7$K and $T_{h}=1$K. (c) For $T_{l}=0.7$K and $T_{h}=1.5$K. (d) For $T_{l}=1$K and $T_{h}=1.5$K.
In Fig.10, we present the quantum operational modes of the Stirling cycle for an engine using a ${\rm Cu}_{3}$-As compound in the $B_{0}-B_{1}$ plane. Panel (a) illustrates this for fixed temperatures $T_{l}=0.5$K and $T_{h}=1$K. It is clear that when $B_{1}>B_{0}$, the system mainly operates as a heat engine. Conversely, when $B_{1}<B_{0}$, the system functions as a refrigerator. Interestingly, around $B_{0}\sim 4.5$T and $B_{1}\sim 3$T, there is a region where it operates as a heat engine. In contrast, around $B_{0}\sim 3$T and $B_{1}\sim 4.5$T, the system functions as an accelerator. There is even a small region at $B_{0}\sim 3$T and $B_{1}\sim 5$T where it operates as a heater. Additionally, between these two main operational modes, there are regions where the heater and thermal accelerator modes occur, particularly when $B_{1}\lesssim B_{0}$. Panel (b) depicts the scenario for $T_{l}=0.7$K and $T_{h}=1$K, showing a similar pattern to panel (a). The main difference is that the heater operation mode is significantly reduced. However, at $B_{0}\sim 3$T and $B_{1}\sim 5$, the system operates as a refrigerator, surrounded by the heater operation mode. Moreover, for $B_{0}\sim 4.5$T and $B_{1}\lesssim 1$T, the system acts as a refrigerator. In Panel (c), we plot the results for $T_{l}=0.7$K and $T_{h}=1.5$K, which is again similar to (b). However, there are no heater or refrigerator modes at $B_{0}\sim 3$T and $B_{1}\sim 5$, though the accelerator mode still remains at $B_{0}\sim 4.5$T and $B_{1}\lesssim 1$T. Finally, Panel (d) presents results for a different set of parameters, $T_{l}=1$K and $T_{h}=1.5$K. Here, we observe wider regions for the heater and accelerator modes, but the inversion operation mode still occurs at $B_{0}\sim 3$T and $B_{1}\sim 4.5$T, and vice versa.
@@
Figure 11: Thermal efficiency of the quantum Stirling cycle for ${\rm Cu}_{3}$-As compound in the $B_{0}$-$B_{1}$ plane. The cycle is evaluated under varying cold ($T_{l}$) and hot ($T_{h}$) temperature conditions: (a) For $T_{l}=0.5$ and $T_{h}=1$. (b) For $T_{l}=0.7$ and $T_{h}=1$. (c) For $T_{l}=0.7$ and $T_{h}=1.5$. (d) For $T_{l}=1$ and $T_{h}=1.5$.